OpenStudy (anonymous):

A toy company hires you to construct a device that uses a small spring to launch 15 g foam pellets. The spring requires an average force of 5.70 N to compress it 1.30 cm. Determine the speed of one of the foam pellets as it leaves the spring

OpenStudy (abhisar):

Do u know how to calculate potential energy of a spring ?

OpenStudy (anonymous):

ep=1/2kx^2?

OpenStudy (abhisar):

yes, now as the spring leaves...this potential energy is transferred to the pellet as its kinetic energy.

OpenStudy (abhisar):

So, \(\huge\tt \frac{1}{2}kx^2=\frac{1}{2}mv^2\)

OpenStudy (anonymous):

do u manipulate the formula for v then?

OpenStudy (abhisar):

yes, u'll have to plug in the values and find v

OpenStudy (abhisar):

Note: value of k is not given

OpenStudy (anonymous):

how do u find k?

OpenStudy (abhisar):

What will you do now ?

OpenStudy (abhisar):

F=kx, so k=F/x Put this in the equation above.

OpenStudy (abhisar):

F and x are given

OpenStudy (anonymous):

x=1.30cm f=5.70N right?

OpenStudy (abhisar):

yes...now just plug in the values and find v

OpenStudy (anonymous):

hi can you please send me the formula again it says error where the p0icture showed up yesterday? @Abhisar

OpenStudy (abhisar):

???

OpenStudy (anonymous):

oh its there again nvm! sorry lol

OpenStudy (abhisar):